An airplane can travel 320 mph in still air. If it travels 690 miles with the wind in the same length of time it travels 510 miles against the wind, what is the speed of the wind? I have the answer I just need help creating an equation for this word problem. The answer is 48 mph.
Answer by scott8148(6628)

t = d / r ___ the times are equal

690 / (320 + w) = 510 / (320 - w)

"cross" multiplying is a good start